OpenStudy (anonymous):

Solve. 1/25 = 5^(x + 4) a. x = –7/2 b. x = –6 c. x = 9/2 d. x = 2

OpenStudy (anonymous):

For this one, you need to convert (1/25) to an exponent with base 5

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

1/25 = 1/(5^2) right?

OpenStudy (anonymous):

When the exponent is also located in the denominator, you can switch the term to the numerator and give the exponent a negative sign. So 1/25 = 1/(5^2) = 5^-2 Try to plug these into a calculator; they all yield the same result.

OpenStudy (anonymous):

I still don't get it. Sorry :/ my teachers don't know how to teach.

OpenStudy (anonymous):

That's ok! :) we're here to help! :)

OpenStudy (anonymous):

A negative exponent is the same as having 1/something to the positive exponent. Ex) 4^-2 is equivalent to 1/(4^2) does that make sense?

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\] \[\frac{1}{25}=\frac{1}{5^2}=5^{-2}\]

OpenStudy (anonymous):

So which answer is it?

OpenStudy (anonymous):

If you can use what we've shown you, we can confirm it :) we're here to help you learn, not just give you the answers. I know it's hard sometimes, but if you learn it now, you won't have to worry about it later :)

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Are you just guessing?

OpenStudy (anonymous):

\[5^{-2}=5^{x+4}\implies -2=x+4\] \[-2=x+4\] \[x=-6\]

OpenStudy (anonymous):

Nope. And thank you!

OpenStudy (anonymous):

*sigh* satellite73 is correct